# ReSpark: Leveraging Previous Data Reports as References to Generate New Reports with LLMs

### v2

@@ Abstract.
Figure 1. ReSpark Overview. ReSpark helps users generate new data reports by reusing existing ones. It begins by ranking reports based on their alignment with the target dataset to assist in selecting a reference report (a). The selected report is then segmented into a sequence of analysis segments (b). For each segment, ReSpark extracts the analysis objective (c) and adapts it to the target dataset (d1). It then generates the corresponding code and charts (d2), followed by textual insights (d3). These elements are finally composed into a new report.
[…]
This figure illustrates the ReSpark workflow. (a) Users first view a dataset and select a reference report ranked by alignment. (b) The selected report is split into analysis segments. (c) Each segment contains an analytical objective, a chart, and a narrative. (d) ReSpark adapts each segment to the target dataset by modifying the objective (d1), generating new code and charts (d2), and rewriting the narrative (d3), which are composed into a new report.
@@ 3.1. Definition of Data Reports
Figure 2. Producing a data report (c1) involves analyzing the data (a) and summarizing the analyzed insights into a data report (b). Specifically, the data analysis workflow (a) includes a series of interdependent analysis segments (a1), each corresponding to an analytical objective, data transformations, and insights (a2). To reuse an existing report on a new dataset, we first deduce its data analysis workflow and reproduce it on the new data (c2).
[…]
This figure illustrates how data reports are generated and reused. (a) The data analysis workflow begins with a dataset (a1) and proceeds through interdependent analysis segments, each consisting of an analytical objective, data transformations, and insights (a2). (b) These insights are organized into a data report. To reuse a report for new data, ReSpark first deduces the original analysis logic (c1) and then reproduces it with the target dataset (c2).
@@ 4.1.2. Report Segmentation
Figure 3. The interface of ReSpark. ReSpark consists of four views: data view (b-c), dependency view (d-e), content view (f-g), and generation view (h-k). The data view displays the dataset description and field information. The dependency view displays the extracted interdependent analysis segments. The content view shows the analytical objective and content of the selected segment. The generation view shows real-time generated results.
[…]
This figure shows the interface of ReSpark, consisting of four main views: (b–c) Data View displays the dataset’s description and detailed data fields. (d–e) Dependency View presents the extracted analysis segments and their dependencies. (f–g) Content View shows the selected segment’s analytical objective, chart, and narrative. (h–k) Generation View supports generating and editing new segments based on the target dataset, including defining new objectives (h), generating code (i), executing results (j), and producing narratives (k). The system enables users to trace, inspect, and adapt analysis logic interactively.
@@ 5. Interface
Figure 4. A demonstration of the ranked reports and their summarized information.
[…]
The figure shows the interface for selecting a reference report in ReSpark. On the left, a ranked list displays available reports along with their titles, topics, predicted data fields, and file sizes. Users can preview report content or remove reports using action buttons. On the right, a preview panel shows the selected report’s title, summary text, and an example chart, helping users assess its relevance before selection.
@@
Figure 5. The process of generating the second segment.
[…]
An illustration of how ReSpark generates the second segment by reusing a reference. (a) shows the second segment from the original Chicago report, analyzing which crime types drove the overall increase. It includes an objective, a textual insight highlighting theft and car theft, and a bar chart. (b) shows the corresponding segment in the generated Los Angeles report, adapted for a different time range and dataset. It reuses the analytical structure but updates the content to reflect new leading crime types such as petty theft.
@@
Figure 6. A demonstration of an analytical objective that fails to be corrected and needs to be removed.
[…]
This figure shows a segment that needs to be removed in ReSpark. (a) The original objective asks about the change in homicides from 2019 to 2022 in Chicago. The rewritten version attempts to generalize it to a longer historical period (since 2000), which the target dataset does not support. (b) ReSpark identifies that the target dataset only spans from 2020 to 2023 and therefore cannot answer the generalized question, marking the objective as invalid and suggesting its removal.
@@
Figure 7. A demonstration of inserting an analysis segment and generating a new analytical objective.
[…]
This figure shows how a user inserts a new analysis segment in ReSpark. The user first adds a new segment and selects a relevant data field (“Time Occ”) and logic pattern (“similarity”). ReSpark then generates a new analytical objective: “What are the trends of arson incidents in Los Angeles based on the time of day?”
@@
Figure 8. A demonstration of generating new titles and adding section structures.
[…]
This figure demonstrates how users can organize the report structure in ReSpark by generating a new report title and grouping related analysis segments into sections. The left side shows the automatic generation of a new report title based on updated content. The right side illustrates how users can add section headers and reorganize segments to improve report clarity.
@@
Figure 9. A demonstration of highlighting the sentences that serve non-data analysis purposes.
[…]
This figure highlights subjective or non-data-driven statements in both a reference COVID-19 report (left) and the generated report (right). Sentences with speculative language or policy commentary are shaded to show their narrative nature. This illustrates how ReSpark can identify and flag subjective content that goes beyond data analysis.
@@ 6.1.4. Procedure and Tasks
Figure 10. Results of the 5-point Likert scale questionnaire in the comparative study. (a) Overall results summarizing all ratings across the three report sets. (b) Detailed results for each report set. Error bars indicate standard deviations (SD).
[…]
A grid of bar charts displays Likert scale ratings from the comparative study evaluating report quality. Subfigure (a) summarizes all user ratings across three report sets. Subfigures (b1), (b2), and (b3) present results for each report set individually. Each chart includes bars representing three versions of reports: baseline (teal), ablation (light green), and ReSpark (dark blue), with error bars showing standard deviations. ReSpark consistently receives higher ratings across most criteria.
@@ 6.2.3. Issues in Chart Designs and Text-Chart Consistency
Figure 11. Example of an inappropriate chart generated in the Titanic dataset report, where the distinction between survivors and non-survivors is made using two colors.
[…]
An example of an inappropriate chart from the Titanic dataset report. Two histograms are overlaid using color to distinguish survivors (blue) and non-survivors (red). The color scheme alone makes it difficult to interpret overlapping data, leading to potential misinterpretation. This highlights the importance of effective visual encoding in automatically generated charts.
@@ 7.1.4. Procedure
Figure 12. User ratings for different features of ReSpark.
[…]
User satisfaction ratings for different features of ReSpark. Each bar represents responses on a 5-point Likert scale, from “Very Dissatisfied” to “Very Satisfied.” Across features such as ranking, segmentation, generation, modification, and time cost, the majority of users reported being satisfied or very satisfied.
Figure 13. Interaction sequences from the usability study, including user information with assigned datasets and selected reference reports with their ranks (a) and interaction logs during reference report selection (b) and new report generation (c).
[…]
Interaction log from the usability study. Panel (a) shows user assignments, selected reference reports, and their ranks. Panel (b) visualizes participants’ interactions during reference report selection. Panel (c) presents the full interaction sequences during report generation, including system-generated and user-modified objectives, content, and segment edits.

### v3

@@ Abstract.
Figure 1. ReSpark Overview. ReSpark helps users generate new data reports by reusing existing ones. It begins by ranking reports based on their alignment with the target dataset to assist in selecting a reference report (a). The selected report is then segmented into a sequence of analysis segments (b). For each segment, ReSpark extracts the analysis objective (c) and adapts it to the target dataset (d1). It then generates the corresponding code and charts (d2), followed by textual insights (d3). These elements are finally composed into a new report.:
@@ 3.1. Definition of Data Reports
Figure 2. Producing a data report (c1) involves analyzing the data (a) and summarizing the analyzed insights into a data report (b). Specifically, the data analysis workflow (a) includes a series of interdependent analysis segments (a1), each corresponding to an analytical objective, data transformations, and insights (a2). To reuse an existing report on a new dataset, we first deduce its data analysis workflow and reproduce it on the new data (c2). :
@@ 4.1.2. Report Segmentation
Figure 3. The interface of ReSpark. ReSpark consists of four views: data view (b-c), dependency view (d-e), content view (f-g), and generation view (h-k). The data view displays the dataset description and field information. The dependency view displays the extracted interdependent analysis segments. The content view shows the analytical objective and content of the selected segment. The generation view shows real-time generated results. :
@@ 5. Interface
Figure 4. A demonstration of the ranked reports and their summarized information. :
@@
Figure 5. The process of generating the second segment. :
@@
Figure 6. A demonstration of an analytical objective that fails to be corrected and needs to be removed. :
@@
Figure 7. A demonstration of inserting an analysis segment and generating a new analytical objective. :
@@
Figure 8. A demonstration of generating new titles and adding section structures. :
@@
Figure 9. A demonstration of highlighting the sentences that serve non-data analysis purposes. :
@@ 6.1.4. Procedure and Tasks
Figure 10. Results of the 5-point Likert scale questionnaire in the comparative study. (a) Overall results summarizing all ratings across the three report sets. (b) Detailed results for each report set. Error bars indicate standard deviations (SD). :
@@ 6.2.3. Issues in Chart Designs and Text-Chart Consistency
Figure 11. Example of an inappropriate chart generated in the Titanic dataset report, where the distinction between survivors and non-survivors is made using two colors. :
@@ 7.1.4. Procedure
Figure 12. User ratings for different features of ReSpark.:
Figure 13. Interaction sequences from the usability study, including user information with assigned datasets and selected reference reports with their ranks (a) and interaction logs during reference report selection (b) and new report generation (c).: